OpenStudy (anonymous):

The difference between two numbers is 5. If the smaller is doubled, the result will be two less than the larger. Find the numbers?? I need HELP please :/

OpenStudy (anonymous):

Lets say x is the bigger number and y is the smaller number x-y=5 2y=x-2 Now simultaneously solve them.

OpenStudy (anonymous):

I am followingg off what williamnz said. Add y to both sides of the first equation to get x by itself and substitute x in the second equation.

OpenStudy (anonymous):

You can either use the substitution or elimination method. I'll show you the elimination one x-y=5 x-2y=2 Eliminate x -y-(-2y)=5-2 y=5-2 y=3 Now to find x we input y into one of the equations x-3=5 x=5+3 x=8 So the numbers are 3, 8

OpenStudy (anonymous):

@williamnz i was actually just going to ask what method do i use but you answer my question thank you

OpenStudy (anonymous):

no problem